### v1

# A hidden symmetry of complex spacetime and the emergence of the standard model algebraic structure

Abstract
When spacetime is considered as a subspace of a wider complex spacetime manifold, there is a mismatch of the elementary linear representations of their symmetry groups, the real and complex Poincaré groups. In particular, no spinors are allowed for the complex case. When a spinh structure is implemented on principal bundles in complex spacetime, one is naturally led to an algebraic structure analogous to the one of the standard model.
1 Introduction: Real slices in complex spacetime
In the past, several authors have studied extensions of physical theories to the complex domain. Einstein in 1945 [1] generalized his gravitational theory to a metric with complex components, others constructed solutions of the complex Einstein equation [2] [3] or a complex manifold description of massless particles [4] [5], etc.
A more radical hypothesis would be to assume that the spacetime manifold itself is parametrized by complex coordinates. Explorations in this direction have been carried out, for example [6] [7], and a more complete set of references and perspectives may be obtained from Ref.[8]. A important question that was left open is, if the spacetime dimensions are indeed indexed by complex numbers, why we do not feel it in our everyday experience. This question has been revisited [9] for spacetimes where coordinates take values in normed division algebras, complex, quaternion or octonion. In these extended spacetimes, any four independent directions may be chosen as a basis for a real vector space, that is, for a real slice of the complex spacetime. Using as consistency condition that the symmetry groups of the extended spacetime must reduce to the real Lorentz and Poincaré groups in each real slice, one obtains, for the homogeneous symmetry group of the extended spacetime, the condition
$\Lambda^{{\dagger}}G\Lambda=G$ (1)
that is, a $U\left(1,3,\mathbb{K}\right)$ group, $\mathbb{K=C},\mathbb{Q},\mathbb{O}$, with $G$ the metric $\left(1,-1,-1,-1\right)$. For $\mathbb{K=C}$, together with the inhomogeneous translations,
[…]
this is the group that, in the past, has been called the Poincaré group with real metric [10] [11] [12]. It is a $16$ parameter complex Lorentz group distinct from the $12$ parameter group
$\Lambda^{T}G\Lambda=G$ (3)
used in the analytic continuations of the S-matrix. Nevertheless it is the condition (1) that insures an identical group in each one of the real slices.
The nature of the matter states, that is implied by this group structure, was obtained [9] by studying the representations of the semidirect group (2). The conclusions were:
- Half-integer spin states are elementary states only for $\mathbb{K=R}$, not for the other normed algebras. Half-integer elementary states, being irreducible representations of the real groups, cannot communicate between different real slices, in the sense that they cannot be rotated from one real slice to another.
- Integer-spin states are elementary states of the higher groups. However these states have a particular nature when the full groups, including reflections, are taken into account. For example, in the complex Lorentz group (1) parity and time reversal are continuously connected to the identity. Therefore, in faithful continuous norm-preserving representations, both parity and time reversal must be implemented by unitary operators. On the other hand one knows that, in each real slice, positivity of energy requires an antiunitary time reversal operator.
In conclusion: In the complex spacetime scenario there are three types of states: half-integer spin states that are confined to the real slices, integer-spin positive energy states that are also only associated to the real slices and finally integer-spin states, with unitary time reversal operator, that can communicate between real slices. This may have physical consequences for time reversal violation and dark matter [13]. The conclusions are qualitatively similar for the normed division algebras $\mathbb{Q}$ and $\mathbb{O}$.
2 Consistency of the symmetry representations and emergence of an algebraic structure
The generators of the Lie algebra of the complex Poincaré group are $\left\{M_{\mu\nu},N_{\mu\nu},K_{\mu},H_{\mu}\right\}$$\left\{M_{\mu\nu},N_{\mu\nu}\right\}$ being the generators of the $U(1,3)$ algebra, $P_{\mu}$ and $H_{\mu}$ the generators of real and complex translations, with commutation relations
$\left[M_{\mu\nu},M_{\rho\sigma}\right]$ $=$ $-M_{\mu\sigma}g_{\nu\rho}-M_{\nu\rho}g_{\mu\sigma}+M_{\nu\sigma}g_{\rho\mu}+M_{\mu\rho}g_{\nu\sigma}$
$\left[M_{\mu\nu},N_{\rho\sigma}\right]$ $=$ $-N_{\mu\sigma}g_{\nu\rho}+N_{\nu\rho}g_{\mu\sigma}+N_{\nu\sigma}g_{\rho\mu}-N_{\mu\rho}g_{\nu\sigma}$
$\left[N_{\mu\nu},N_{\rho\sigma}\right]$ $=$ $M_{\mu\sigma}g_{\nu\rho}+M_{\nu\rho}g_{\mu\sigma}+M_{\nu\sigma}g_{\rho\mu}+M_{\mu\rho}g_{\nu\sigma}$
$\left[M_{\mu\nu},P_{\rho}\right]$ $=$ $-g_{\nu\rho}P_{\mu}+g_{\mu\rho}P_{\nu}$
$\left[M_{\mu\nu},H_{\rho}\right]$ $=$ $-g_{\nu\rho}H_{\mu}+g_{\mu\rho}H_{\nu}$
$\left[N_{\mu\nu},P_{\rho}\right]$ $=$ $-g_{\nu\rho}H_{\mu}-g_{\mu\rho}H_{\nu}$
$\left[N_{\mu\nu},H_{\rho}\right]$ $=$ $g_{\nu\rho}P_{\mu}+g_{\mu\rho}P_{\nu}$
$\left[P_{\mu},H_{\rho}\right]$ $=$ 00 0 (4)
Several authors in the past, looking for spectrum generating algebras, have dismissed the $U(1,3)$ algebra, as useless, for not having irreducible spinor representations. However, the fact that $O(1,3)$ has spinor representations and $U(1,3)$ does not, is the most interesting feature of the latter.
If however $U(1,3)$ is an invariance of Nature, in a complex spacetime context, the apparent spectrum mismatch of the symmetries in the full space with those of the real slices has interesting consequences. The fact that there are no linear irreducible spinor representations, does not mean that such structures might not be present in a nonlinear realization.
Here I feel convenient to analyse the mismatch for the massive representations of the complex Poincaré group, that is, for non-zero eigenvectors of the $P_{\mu}P^{\mu}+H_{\mu}H^{\mu}$ operator. The isotropy group [9] is $U(3)$ which one splits as $U(1)\oplus SU(3)$. Denote the generators of $U(3)$ as $\left\{\mathcal{G}_{i};i=1,\cdots,9\right\}=\left\{R_{1},R_{2},R_{3},U_{1},U_{2},U_{3},D_{1},D_{2},D_{3}\right\}$ [^1] A representation in terms of the usual $SU\left(3\right)$ $\lambda$-matrices is: $R_{1}\rightarrow i\lambda_{7};\;R_{2}\rightarrow-i\lambda_{5};\;R_{3}\rightarrow i\lambda_{2};U_{1}\rightarrow i\lambda_{6};U_{2}\rightarrow i\lambda_{4};U_{3}\rightarrow i\lambda_{1};D_{1}\rightarrow i\lambda_{3};D_{2}\rightarrow\frac{i}{2}\lambda_{8};D_{3}\rightarrow i\sqrt{\frac{2}{3}}\mathds{1}$
$R_{i}$ $=$ $\frac{1}{2}\epsilon_{ijk}M_{jk};U_{i}=\frac{1}{2}\epsilon_{ijk}N_{jk};D_{1}=\frac{1}{2}\left(N_{11}-N_{22}\right);$
$D_{2}$ $=$ $\frac{1}{2\sqrt{3}}\left(N_{11}+N_{22}-2N_{33}\right);D_{3}=\frac{1}{\sqrt{6}}\left(N_{11}+N_{22}+N_{33}\right)$ (5)
the first eigth being the generators of $SU\left(3\right)$ and $\mathcal{G}_{9}$ the generator of $U\left(1\right)$. In a standard $3\times 3$ representation they are trace orthogonal
$Tr\left(\mathcal{G}_{i}\mathcal{G}_{j}\right)=-2\delta_{ij}$ (6)
With a coset decomposition, $SU(3)$ becomes a $SO(3)$ principal bundle $\mathcal{C}\left(\mathcal{C}_{b},\mathcal{C}_{f},SO(3)\right)$, with base the Riemmanian manifold $\mathcal{C}_{b}=SU(3)/SO(3)$ and fibers $\mathcal{C}_{f}\simeq SO(3)$. The metric in $\mathcal{C}_{b}$ is obtained from (6). A coordinate system in $\mathcal{C}$ is obtained by choosing a point in each coset
$\left(\omega^{i}\right)=\exp\left(\omega^{i}\mathcal{G}_{i};i=4,\cdots,8\right)$
and when an element $g\in SU(3)$ acts upon $\left(\omega^{i}\right)$
$g\left(\omega^{i}\right)=\left(\omega^{i\prime}\right)h\left(\omega^{i},g\right)$
with $h\left(\omega^{i},g\right)\in SO\left(3\right)$.
To define a spin structure on $\mathcal{C}$ means to construct a principal fiber bundle over $\mathcal{C}_{b}$ with typical fiber $Spin(3)$ , such that the image of its transition functions [14] are the transition functions of $SO\left(3\right)$. It is known that a $SO\left(n\right)$ bundle over some base admits a spin-bundle extension if and only if the second Stiefel-Whitney homology class, with $\mathbf{Z}_{2}$ coefficients, $w_{2}$ is trivial. However
$H^{2}\left(\mathcal{C}_{b};\mathbf{Z}_{2}\right)=\mathbf{Z}_{2},$
therefore $w_{2}\left(\mathcal{C}_{b}\right)\neq 0$ and the $SO\left(3\right)$ bundle $\mathcal{C}$, by itself, cannot carry a spin structure. This is a known fact because $SU(3)/SO(3)$ is the Wu manifold [15] [16] which is not a $Spin$ manifold, nor a $Spin^{c}$ manifold. However it is a $Spin^{h}$ manifold [17]. In constructive terms what this means is that to implement a spin structure in the $SO(3)$ principal bundle over $\mathcal{C}_{b}$, one should replace $\mathcal{C}$ by a Whitney sum
[…]
with the second bundle $\mathcal{F}$, over the same base $\mathcal{C}_{b}$, carrying a $SU(2)$ structure, the fibers being the direct sum of the fibers of the two bundles.
This has a simple interpretation in terms of the representations of the transition functions. As stated before, the construction of the spin structure implies the establishment of a matching of the transition functions of the $SO(3)$ bundle with those of the corresponding spin bundle. The transition functions are elements of the structure groups and therefore the mathcing should occur at the level of its representations. The set $\left\{R_{i}\right\}$ of $SO\left(3\right)$ generators carries a spin one representation, which does not match the spin $1/2$ of the spin structure. Therefore one needs an additional $SU\left(2\right)$ spin $1/2$, which, coupled with the first one, matches the spin one of the $SO\left(3\right)$ bundle.
[…]
Notice also that this $U\left(1\right)\oplus SU\left(3\right)\oplus SU\left(2\right)$ algebraic structure is not exactly the one of the standard model. In particular the $SU(3)$ is not a purely internal symmetry for it contains the generators of the $O(3)$ group. That implies that an action of $SU(3)$ not only changes the fiber (colour?) but also spin, which might have observable consequences.
Acknowledgments
Partially supported by Fundação para a Ciência e a Tecnologia (FCT), project UIDB/04561/2020: https://doi.org/10.54499/UIDB/04561/2020