### v2

# A hidden symmetry of complex spacetime and the emergence of the standard model algebraic structure

Abstract
When spacetime is considered as a subspace of a wider complex spacetime manifold, there is a mismatch of the elementary linear representations of their symmetry groups, the real and complex Poincaré groups. In particular, no spinors are allowed for the complex case. When a spinh structure is implemented on principal bundles in complex spacetime, one is naturally led to an algebraic structure similar to the one of the standard model.
1 Introduction: Real fibers in complex spacetime
In the past, several authors have studied extensions of physical theories to the complex domain. Einstein in 1945 [1] generalized his gravitational theory to a metric with complex components, others constructed solutions of the complex Einstein equation [2] [3] or a complex manifold description of massless particles [4] [5], etc.
A more radical hypothesis would be to assume that the spacetime manifold itself is parametrized by complex coordinates. Explorations in this direction have been carried out, for example [6] [7], and a more complete set of references and perspectives may be obtained from Ref.[8]. A important question that was left open is, if the spacetime dimensions are indeed indexed by complex numbers, why we do not feel it in our everyday experience. This question has been revisited for spacetimes where coordinates take values in normed division algebras, complex, quaternion or octonion [9]. For the complex field case, a complex manifold $\mathcal{M}$ is considered, with (local) metric $G=\left(1,-1,-1,-1\right)$ and symmetry group $L_{\mathbb{C}}\mathcal{=}U\left(1,3\right)$
$\Lambda^{{\dagger}}G\Lambda=G$ (1)
$\left(\Lambda\in L_{\mathbb{C}}\right)$, the real spacetime being considered to be a $4-$dimensional subspace $\mathcal{S}$ with Lorentzian metric, that is, a Lorentzian fiber in the Grassmannian $Gr_{4}\left(\mathcal{M}\right)$ of four-dimensional frames of $\mathcal{M}$. The symmetry group of $\mathcal{S}$ is the real Lorentz group $L_{\mathbb{R}}\mathcal{=}SO\left(1,3\right)$. Together with the complex inhomogeneous translations, the (local) symmetry group of $\mathcal{M}$ is
[…]
$P_{\mathbb{C}}$ is the group that, in the past, has been called the Poincaré group with real metric [10] [11] [12]. It is a $24-$parameter complex group distinct from the $16-$parameter group used in the analytic continuations of the S-matrix.
The nature of the matter states, that is implied by this group structure, was obtained [9] by studying the representations of the semidirect group (2). The conclusions were:
- Half-integer spin states are elementary states only for the real Lorentzian fibers, not for the complex manifold $\mathcal{M}$, nor for the higher normed algebras. Their absence as elementary states in the complex manifold $\mathcal{M}$ may be physically interpreted as meaning that such states cannot be ”rotated” away from each real Lorentzian fiber to the other fibers of the Grassmannian $Gr_{4}\left(\mathcal{M}\right)$.
- Integer-spin states are elementary states of the complex group. However a superselection rule emerges when discrete symmetries are included [13].
Whatever the point of view concerning the extension to complex spacetime, a relevant question is: When real spacetime is embedded into a higher dimensional division algebra structure, what properties of the higher structure are inherited or non-inherited by the real spacetime? The absence of spinor states for the complex $P_{\mathbb{C}}$ group is particularly interesting. It might be that $P_{\mathbb{C}}$ is actually also the symmetry group of the real Lorentzian fiber, but that the symmetry is implemented in a subtler way. This second point of view leads to an analysis of when spin structures are or are not realized in coset manifolds. And to what additional quantum numbers appear when one insists on their implementation.
2 Consistency of the symmetry representations and emergence of an algebraic structure
The generators of the Lie algebra of the complex Poincaré group $P_{\mathbb{C}}$ are $\left\{M_{\mu\nu},N_{\mu\nu},K_{\mu},H_{\mu}\right\}$ , $M_{\mu\nu}$ and $N_{\mu\nu}$ being the generators of the $U(1,3)$ algebra, $P_{\mu}$ and $H_{\mu}$ the generators of real and complex translations, with commutation relations
$\left[M_{\mu\nu},M_{\rho\sigma}\right]$ $=$ $-M_{\mu\sigma}g_{\nu\rho}-M_{\nu\rho}g_{\mu\sigma}+M_{\nu\sigma}g_{\rho\mu}+M_{\mu\rho}g_{\nu\sigma}$
$\left[M_{\mu\nu},N_{\rho\sigma}\right]$ $=$ $-N_{\mu\sigma}g_{\nu\rho}+N_{\nu\rho}g_{\mu\sigma}+N_{\nu\sigma}g_{\rho\mu}-N_{\mu\rho}g_{\nu\sigma}$
$\left[N_{\mu\nu},N_{\rho\sigma}\right]$ $=$ $M_{\mu\sigma}g_{\nu\rho}+M_{\nu\rho}g_{\mu\sigma}+M_{\nu\sigma}g_{\rho\mu}+M_{\mu\rho}g_{\nu\sigma}$
$\left[M_{\mu\nu},P_{\rho}\right]$ $=$ $-g_{\nu\rho}P_{\mu}+g_{\mu\rho}P_{\nu}$
$\left[M_{\mu\nu},H_{\rho}\right]$ $=$ $-g_{\nu\rho}H_{\mu}+g_{\mu\rho}H_{\nu}$
$\left[N_{\mu\nu},P_{\rho}\right]$ $=$ $-g_{\nu\rho}H_{\mu}-g_{\mu\rho}H_{\nu}$
$\left[N_{\mu\nu},H_{\rho}\right]$ $=$ $g_{\nu\rho}P_{\mu}+g_{\mu\rho}P_{\nu}$
$\left[P_{\mu},H_{\rho}\right]$ $=$ 00 0 (3)
Several authors in the past, looking for spectrum generating algebras, have dismissed the $U(1,3)$ algebra, as useless, for not having irreducible spinor representations. However, the fact that $O(1,3)$ has spinor representations and $U(1,3)$ does not, is the most interesting feature of the latter.
If however $U(1,3)$ is an invariance of Nature, in a complex spacetime context, the apparent spectrum mismatch of the symmetries in the full space with those of the real Lorentzian fibers has interesting consequences. The fact that there are no linear irreducible spinor representations, does not mean that such structures might not be present in a nonlinear realization.
Let us analyze the mismatch for the massive representations of the complex Poincaré group $P_{\mathbb{C}}$, that is, for non-zero eigenvectors of the $P^{2}=P_{\mu}P^{\mu}+H_{\mu}H^{\mu}$ operator. The isotropy group [9] is $U(3)$ which one splits as $U(1)\oplus SU(3)$. Denote the generators of $U(3)$ as $\left\{\mathcal{G}_{i};i=1,\cdots,9\right\}=\left\{R_{1},R_{2},R_{3},U_{1},U_{2},U_{3},D_{1},D_{2},D_{3}\right\}$ [^1] A representation in terms of the usual $SU\left(3\right)$ $\lambda$-matrices is: $R_{1}\rightarrow i\lambda_{7};\;R_{2}\rightarrow-i\lambda_{5};\;R_{3}\rightarrow i\lambda_{2};U_{1}\rightarrow i\lambda_{6};U_{2}\rightarrow i\lambda_{4};U_{3}\rightarrow i\lambda_{1};D_{1}\rightarrow i\lambda_{3};D_{2}\rightarrow\frac{i}{2}\lambda_{8};D_{3}\rightarrow i\sqrt{\frac{2}{3}}\mathds{1}$
$R_{i}$ $=$ $\frac{1}{2}\epsilon_{ijk}M_{jk};U_{i}=\frac{1}{2}\epsilon_{ijk}N_{jk};D_{1}=\frac{1}{2}\left(N_{11}-N_{22}\right);$
$D_{2}$ $=$ $\frac{1}{2\sqrt{3}}\left(N_{11}+N_{22}-2N_{33}\right);D_{3}=\frac{1}{\sqrt{6}}\left(N_{11}+N_{22}+N_{33}\right)$ (4)
the first eight being the generators of $SU\left(3\right)$ and $\mathcal{G}_{9}$ the generator of $U\left(1\right)$. In a standard $3\times 3$ representation they are trace orthogonal
$Tr\left(\mathcal{G}_{i}\mathcal{G}_{j}\right)=-2\delta_{ij}$ (5)
With a coset decomposition, $SU(3)$ becomes a $SO(3)$ principal bundle $\mathcal{C}\left(\mathcal{C}_{b},\mathcal{C}_{f},SO(3)\right)$, with base the Riemmanian manifold $\mathcal{C}_{b}=SU(3)/SO(3)$ and fibers $\mathcal{C}_{f}\simeq SO(3)$. The metric in $\mathcal{C}_{b}$ is obtained from (5). A coordinate system in $\mathcal{C}$ is obtained by choosing a point in each coset
$\left(\omega^{i}\right)=\exp\left(\omega^{i}\mathcal{G}_{i};i=4,\cdots,8\right)$ (6)
and when an element $g\in SU(3)$ acts upon $\left(\omega^{i}\right)$
$g\left(\omega^{i}\right)=\left(\omega^{i\prime}\right)h\left(\omega^{i},g\right)$ (7)
with $h\left(\omega^{i},g\right)\in SO\left(3\right)$.
To define a spin structure on $\mathcal{C}$ means to construct a principal fiber bundle over $\mathcal{C}_{b}$ with typical fiber $Spin(3)$ (a double cover of $\mathcal{C}$), such that the image of its transition functions [14] are the transition functions of $SO\left(3\right)$. It is known that a $SO\left(n\right)$ bundle over some base admits a spin-bundle extension if and only if the second Stiefel-Whitney homology class, with $\mathbf{Z}_{2}$ coefficients, $w_{2}$ is trivial. However $SU(3)/SO(3)$ is the Wu manifold which is known [15] [16] not to be a $Spin$ manifold, nor a $Spin^{c}$ manifold. It is however a $Spin^{h}$ manifold [17]. What this means is that to implement a spin structure in the $SO(3)$ principal bundle over $\mathcal{C}_{b}$, one should complement $\mathcal{C}$ by a Whitney sum with a second bundle $\mathcal{P}_{SU\left(2\right)}^{\prime}$, over the same base $\mathcal{C}_{b}$, carrying a quaternionic ($SU(2)$) structure, the fibers being the direct sum of the fibers of the two bundles.
This has a simple interpretation in terms of the representations of the transition functions. As stated before, the construction of the spin structure implies the establishment of a matching of the transition functions of the $SO(3)$ bundle with those of the corresponding spin bundle. The transition functions are elements of the structure groups and therefore the matching should occur at the level of its representations. The set $\left\{R_{i}\right\}$ of $SO\left(3\right)$ generators, as a subgroup of $SU\left(3\right)$, carries a spin one representation, which does not match the spin $1/2$ of the spin structure. Therefore one needs an additional $SU\left(2\right)$ spin $1/2$, which, coupled with the first one, matches the spin one of the $SO\left(3\right)$ bundle in the $SU\left(3\right)$ triplet.
Use a creation operator representation of the Lie algebra of $SU\left(3\right)$. Let $\left\{a_{i}^{{\dagger}},a_{i};i=1,2,3\right\}$ be a set of bosonic creation and annihilation operators. Then the generators of the $SU\left(3\right)$ algebra have the representation
$R_{i}=\epsilon_{ijk}a_{j}^{{\dagger}}a_{k};\;U_{i}=i\left|\epsilon_{ijk}\right|a_{j}^{{\dagger}}a_{k};\;D_{1}=i\left(a_{1}^{{\dagger}}a_{1}-a_{2}^{{\dagger}}a_{2}\right);\;D_{2}=\frac{i}{\sqrt{3}}\left(a_{1}^{{\dagger}}a_{1}+a_{2}^{{\dagger}}a_{2}-2a_{3}^{{\dagger}}a_{3}\right),$ (8)
a $SU\left(3\right)$ triplet being formed by $3$ spin-one states
$\left|1\right)=\frac{1}{\sqrt{2}}\left(a_{1}^{{\dagger}}+ia_{2}^{{\dagger}}\right)\left|0\right\rangle;\;\left|-1\right)=\frac{1}{\sqrt{2}}\left(a_{1}^{{\dagger}}-ia_{2}^{{\dagger}}\right)\left|0\right\rangle;\;\left|0\right)=a_{3}^{{\dagger}}\left|0\right\rangle$ (9)
Now one has to match these states by composing an observable spin $\frac{1}{2}$ state in the real Lorentzian fiber with another spin $\frac{1}{2}$ state of the auxiliary $\mathcal{P}_{SU\left(2\right)}^{\prime}$ bundle. Notice however that the three states in (9) not only transform under the $R_{i}$ rotation generators, but also have nontrivial transformation properties under the other $SU\left(3\right)$ generators, in particular $U_{1}$ and $U_{2}$. This must  also be matched by the composite states.
Let $c_{i}^{\uparrow{\dagger}}$ and $c_{j}^{\downarrow{\dagger}}$ be the spin up and spin down creator operators of the observable state in the real Lorentzian fiber. And $b_{i}^{\uparrow{\dagger}}$ and $b_{j}^{\downarrow{\dagger}}$ the corresponding operators of the auxiliary $\mathcal{P}_{SU\left(2\right)}^{\prime}$ bundle. The $i$ and $j$ labels stand for the additional quantum numbers needed to match the $SU\left(3\right)$ transformation properties of the states. The following scheme satisfies the matching requirements
$\left|1\right)\rightarrow b_{i}^{\uparrow{\dagger}}c_{i}^{\uparrow{\dagger}}\left|0\right\rangle;\;\left|-1\right)\rightarrow b_{j}^{\downarrow{\dagger}}c_{j}^{\downarrow{\dagger}}\left|0\right\rangle;\;\left|0\right)\rightarrow\left(b_{i}^{\uparrow{\dagger}}c_{j}^{\downarrow{\dagger}}+b_{j}^{\downarrow{\dagger}}c_{i}^{\uparrow{\dagger}}\right)\left|0\right\rangle$ (10)
Notice that the ”internal” quantum numbers $i$ and $j$ of the spinor states $c_{i}^{\uparrow{\dagger}}$ and $c_{j}^{\downarrow{\dagger}}$ may be different for different spin projections, showing the chiral nature of this quantum number. The spinor doublets in the real Lorentzian fiber are thus endowed with a doublet of ”internal” quantum numbers. This might superficially look like the emergence of an additional $SU\left(2\right)$ structure. However these quantum numbers are matching a structure acted upon by the $U_{i}$ operators which are not a closed $SU\left(2\right)$ structure, but $SU\left(3\right)$ operators. Therefore this emergent structure that will tentatively be called ”flavor” is not $SU\left(2\right)$ but a doublet of ”flavor $SU\left(3\right)$”. This $SU\left(3\right)$ structure, that at the complex spacetime level is an exact symmetry, may connect states with different $p^{2}=P^{\mu}P_{\mu}$values (the real mass square) because they are eigenvalues of $P^{2}=P_{\mu}P^{\mu}+H_{\mu}H^{\mu}$, not of $p^{2}$ and from the commutation relations in (3) one sees how the $U_{i}$ operators permute the $P_{\mu}$ and $H_{\mu}$ operators. Summarizing: an exact complex symmetry may appear as a broken symmetry in the real Lorentzian fibers. An additional feature of the (10) structure is the existence of a spin zero state
$\left(b_{i}^{\uparrow{\dagger}}c_{j}^{\downarrow{\dagger}}-b_{j}^{\downarrow{\dagger}}c_{i}^{\uparrow{\dagger}}\right)\left|0\right\rangle$
With the additional $\mathcal{P}_{SU\left(2\right)}^{\prime}$ bundle, spinor states are consistently constructed at each point of the coset space $SU(3)/SO(3)$. Therefore a second consequence is a degeneracy of the spinor states parametrized by the points of the coset space. This is a $5-$dimensional space [^2] Or $6-$dimensional if instead of $SU(3)/SO(3)$ one consider $U(3)/SO(3)$.. A detailed characterization of this ”color” space and its coordinates (6) is important, particularly if the $SU(3)$ symmetry, in its double role, is gauged.
Summarizing: If the real Poincaré group is a symmetry of the real Lorentzian fibers and $P_{\mathbb{C}}$ a symmetry of the full complex spacetime, consistency of the representations generates a $U\left(1\right)\oplus SU\left(3\right)$ structure with $SU\left(3\right)$ acting in a double role, both as generating additional doublet quantum numbers as a broken $P_{\mu}P^{\mu}$ symmetry and generating a degeneracy of massive spinor states parametrized by the base manifold.
Acknowledgments
Partially supported by Fundação para a Ciência e a Tecnologia (FCT), project UIDB/04561/2020: https://doi.org/10.54499/UIDB/04561/2020